# DepthSync: Diffusion Guidance-Based Depth Synchronization for Scale- and Geometry-Consistent Video Depth Estimation

### v1

@@ Abstract
Diffusion-based video depth estimation methods have achieved remarkable success with strong generalization ability. However, predicting depth for long videos remains challenging. Existing methods typically split videos into overlapping sliding windows, leading to accumulated scale discrepancies across different windows, particularly as the number of windows increases. Additionally, these methods rely solely on 2D diffusion priors, overlooking the inherent 3D geometric structure of video depths, which results in geometrically inconsistent predictions. In this paper, we propose DepthSync, a novel, training-free framework using diffusion guidance to achieve scale- and geometry-consistent depth predictions for long videos. Specifically, we introduce scale guidance to synchronize the depth scale across windows and geometry guidance to enforce geometric alignment within windows based on the inherent 3D constraints in video depths. These two terms work synergistically, steering the denoising process toward consistent depth predictions. Experiments on various datasets validate the effectiveness of our method in producing depth estimates with improved scale and geometry consistency, particularly for long videos.
Figure 1: We present DepthSync, a framework that introduces cross-window depth scale synchronization and intra-window geometry alignment to achieve long video depth predictions with enhanced scale and geometry consistency via diffusion guidance.
@@ 2 Related Work
Monocular Depth Estimation. Monocular depth estimation employs deep neural networks to regress dense depth maps from images. Research progress has been made in both discriminative [11, 14, 31, 56, 2, 3, 21] and generative methods [24, 39] in this field. Notable discriminative methods include MiDaS [38] and Depth Anything [51], which are trained on mixed dataset to enhance generality. Recent progress in generative methods, such as Marigold [28] and GeoWizard [15], demonstrates that the use of diffusion priors can facilitate zero-shot depth estimation with improved prediction details and smoothness. However, despite their impressive performance in complex scenarios, these single-image depth estimation methods often face challenges in maintaining consistency across multiple frames.
@@ 3.1.1 Diffusion Guided Sampling
$\hat{\epsilon}_{\theta}(z_{t},t)=\epsilon_{\theta}(z_{t},t)+s(t)\cdot\nabla_{z_{t}}\mathcal{L}(c,f(\hat{z_{0}}))$ (2)
@@ 4 Experiments
Method Type Video ScanNet GMU KITCHEN Video Bonn KITTI Video ScanNet
@@
DepthAnythingV2 S 150 0.148 0.776 0.188 0.645 190 0.102 0.916 0.151 0.789 150 0.040
Marigold S 0.179 0.740 0.214 0.584 0.098 0.932 0.143 0.820 0.131
NVDS V 0.199 0.647 0.231 0.542 0.163 0.771 0.212 0.658 0.047
ChronoDepth V 0.172 0.749 0.196 0.650 0.092 0.932 0.178 0.733 0.028
DepthCrafter V 0.141 0.799 0.143 0.795 0.095 0.917 0.076 0.967 0.024
DepthSync(Ours) V 0.113 0.870 0.113 0.881 0.069 0.972 0.064 0.980 0.019
DepthAnythingV2 S 210 0.160 0.743 0.189 0.652 270 0.117 0.887 0.159 0.770 210 0.040
Marigold S 0.191 0.710 0.221 0.598 0.113 0.901 0.146 0.809 0.143
NVDS V 0.208 0.622 0.223 0.578 0.176 0.741 0.217 0.647 0.048
ChronoDepth V 0.182 0.726 0.201 0.659 0.112 0.898 0.185 0.707 0.028
DepthCrafter V 0.156 0.757 0.155 0.761 0.121 0.863 0.081 0.960 0.025
DepthSync(Ours) V 0.127 0.836 0.113 0.884 0.077 0.960 0.067 0.977 0.021
DepthAnythingV2 S 270 0.161 0.738 0.181 0.681 350 0.125 0.870 0.163 0.756 270 0.041
Marigold S 0.195 0.699 0.219 0.613 0.123 0.884 0.151 0.796 0.148
NVDS V 0.211 0.614 0.222 0.586 0.180 0.722 0.219 0.643 0.048
ChronoDepth V 0.184 0.716 0.201 0.659 0.124 0.875 0.185 0.705 0.029
DepthCrafter V 0.161 0.745 0.154 0.764 0.130 0.841 0.086 0.951 0.025
DepthSync(Ours) V 0.136 0.814 0.120 0.860 0.083 0.946 0.069 0.975 0.021
DepthAnythingV2 S 330 0.164 0.729 0.191 0.655 430 0.127 0.861 0.161 0.759 330 0.041
Marigold S 0.198 0.691 0.228 0.611 0.126 0.871 0.150 0.800 0.166
NVDS V 0.214 0.608 0.229 0.568 0.182 0.712 0.221 0.641 0.049
ChronoDepth V 0.187 0.709 0.217 0.621 0.130 0.855 0.186 0.706 0.029
DepthCrafter V 0.168 0.725 0.153 0.756 0.134 0.834 0.088 0.949 0.025
DepthSync(Ours) V 0.138 0.802 0.128 0.842 0.090 0.937 0.071 0.974 0.021
DepthAnythingV2 S 390 0.167 0.724 0.193 0.645 510 0.128 0.861 0.159 0.764 390 0.041
Marigold S 0.202 0.689 0.228 0.613 0.126 0.871 0.149 0.800 0.148
NVDS V 0.216 0.603 0.232 0.561 0.185 0.701 0.221 0.641 0.050
ChronoDepth V 0.190 0.704 0.222 0.619 0.132 0.851 0.186 0.705 0.030
DepthCrafter V 0.169 0.721 0.161 0.746 0.138 0.817 0.089 0.947 0.025
DepthSync(Ours) V 0.150 0.780 0.128 0.848 0.091 0.934 0.071 0.973 0.021
DepthAnythingV2 S 450 0.169 0.722 0.193 0.644 590 0.131 0.856 0.159 0.765 450 0.041
Marigold S 0.205 0.686 0.241 0.597 0.128 0.865 0.148 0.803 0.147
NVDS V 0.219 0.597 0.235 0.558 0.189 0.692 0.223 0.640 0.050
ChronoDepth V 0.193 0.699 0.225 0.612 0.132 0.853 0.193 0.691 0.030
DepthCrafter V 0.171 0.716 0.160 0.748 0.143 0.801 0.088 0.948 0.025
DepthSync(Ours) V 0.154 0.769 0.131 0.837 0.093 0.929 0.072 0.972 0.021
@@
150 ATE (m)$\downarrow$ 0.604 0.570 0.522 0.154 0.144
RPE trans (m)$\downarrow$ 0.078 0.132 0.176 0.021 0.020
RPE rot  (deg)$\downarrow$ 2.91 3.41 26.9 0.622 0.611
210 ATE (m)$\downarrow$ 0.741 0.737 0.666 0.211 0.197
RPE trans (m)$\downarrow$ 0.093 0.149 0.197 0.022 0.020
RPE rot  (deg)$\downarrow$ 2.96 4.09 28.3 0.624 0.620
270 ATE (m)$\downarrow$ 0.846 0.854 0.792 0.256 0.242
RPE trans (m)$\downarrow$ 0.098 0.156 0.198 0.023 0.021
RPE rot  (deg)$\downarrow$ 2.96 4.11 29.4 0.641 0.635
330 ATE (m)$\downarrow$ 0.899 0.906 0.849 0.292 0.275
RPE trans (m)$\downarrow$ 0.093 0.158 0.200 0.023 0.021
RPE rot  (deg)$\downarrow$ 2.97 4.11 30.3 0.642 0.627
390 ATE (m)$\downarrow$ 0.941 0.954 0.891 0.334 0.316
RPE trans (m)$\downarrow$ 0.092 0.152 0.204 0.023 0.021
RPE rot  (deg)$\downarrow$ 2.99 4.12 30.5 0.650 0.643
450 ATE (m)$\downarrow$ 1.00 1.01 0.942 0.367 0.350
RPE trans (m)$\downarrow$ 0.094 0.162 0.209 0.023 0.021
RPE rot  (deg)$\downarrow$ 3.12 4.29 30.7 0.676 0.669
@@ 4.2 Zero-Shot Depth Estimation Results
Figure 4: Qualitative Comparison for Long Video Depth Estimation. We compare predictions on 450-frame ScanNet videos and 590-frame Bonn videos. Slicing along the timeline at the green-line position (first column), we concatenate results to visualize temporal changes in color and depth. Red boxes highlight depth scale inconsistencies, while blue boxes mark depth inaccuracies in previous methods. Our method shows superior depth accuracy and scale consistency over time.
@@ Appendix A Implementation Details
A.1 Loss Terms in Geometry Guidance
@@
A.1.1 Global Geometric Constraints constraints
@@
A.1.2 Local Geometric Constraints
@@ Appendix B Comparison with Reconstruction Methods
We compare our method with representative reconstruction methods on the first 10 ScanNet test scenes. COLMAP[40] and 2DGS [23] (uses COLMAP poses) fail on 3 scenes due to failed COLMAP pose estimation, while our method reconstructs all scenes successfully. For shared scenes, our approach achieves more accurate depth and poses (Table 4).
@@ Appendix C Inference Cost and Supplementary Evaluation
Table 5: Inference Cost Evaluation for 90 frames (Resolution: 640$\times$ 448) on a 40GB A100 GPU.
Dataset ScanNet GMU Kitchen Bonn KITTI
@@
ChronoDepth [41] 0.172 0.749 0.196 0.650 0.092 0.932 0.178 0.733
DepthCrafter [22] 0.141 0.799 0.143 0.795 0.095 0.917 0.076 0.967
DepthAnyVideo [50]* 0.100 0.910 0.089 0.932 0.070 0.966 0.085 0.942
DUSt3R [46] 0.059 0.972 0.069 0.968 0.075 0.935 0.155 0.777
MonST3R [59] 0.067 0.959 0.087 0.920 0.046 0.975 0.192 0.692
Ours-S 0.128 0.834 0.126 0.844 0.070 0.968 0.074 0.972
Ours 0.113 0.870 0.113 0.881 0.069 0.972 0.064 0.980
ChronoDepth [41] 0.193 0.699 0.225 0.612 0.132 0.853 0.193 0.691
DepthCrafter [22] 0.171 0.716 0.160 0.748 0.143 0.801 0.088 0.948
DepthAnyVideo [50]* 0.127 0.840 0.113 0.883 0.103 0.911 0.094 0.922
DUSt3R [46] 0.090 0.934 0.127 0.828 0.115 0.889 0.174 0.731
MonST3R [59] 0.099 0.911 0.140 0.817 0.086 0.918 0.232 0.620
Ours-S 0.157 0.757 0.138 0.807 0.094 0.926 0.078 0.967
Ours 0.154 0.769 0.131 0.837 0.093 0.929 0.072 0.972
@@
Comparison with Depth Any Video (DAV) [50]. DAV supports at most 192 frames per video, so we compare with its original implementation only on our shortest video setting (Table 6 ①). For longer videos, we apply the same sliding window strategy as our method, running DAV on each segment and aligning the results. DAV demonstrates degraded performance on long sequences (②) compared to shorter ones (①) over four datasets. Our method outperforms DAV on Bonn and KITTI (Table 6), while DAV performs better on ScanNet and GMU Kitchen - likely due to its extensive training on indoor data (such as Hypersim). In contrast, our base model DepthCrafter has limited exposure to indoor scenes, which may constrain its performance in such environments.
@@ Appendix E Supplementary Qualitative Examples
We provide additional qualitative examples to supplement the main paper in Figure 6, Figure 7, Figure 8 and Figure 9, demonstrating that our DepthSync achieves better scale synchronization between windows for long video and more accurate geometric consistency within each window.

### v2

@@ Abstract
Diffusion-based video depth estimation methods have achieved remarkable success. However, predicting depth for long videos remains challenging. Existing methods typically split videos into overlapping sliding windows, leading to accumulated scale discrepancies across different windows, particularly as the number of windows increases. Additionally, these methods rely solely on 2D diffusion priors, overlooking the inherent 3D geometric structure of video depths, which results in geometrically inconsistent predictions. In this paper, we propose DepthSync, a novel, training-free framework using diffusion guidance to achieve scale- and geometry-consistent depth predictions for long videos. Specifically, we introduce scale guidance to synchronize the depth scale across windows and geometry guidance to enforce geometric alignment within windows based on the inherent 3D constraints in video depths. These two terms work synergistically, steering the denoising process toward consistent depth predictions. Experiments on various datasets validate the effectiveness of our method in producing depth estimates with improved scale and geometry consistency, particularly for long videos.
Figure 1: We present DepthSync, a framework that introduces cross-window depth scale synchronization and intra-window geometry alignment for long video depth predictions with enhanced scale and geometry consistency via diffusion guidance.
@@ 2 Related Work
Monocular Depth Estimation. Monocular depth estimation employs deep neural networks to regress dense depth maps from images. Research progress has been made in both discriminative [11, 14, 31, 57, 2, 3, 21, 38] and generative methods [24, 40] in this field. Notable discriminative methods include MiDaS [39] and Depth Anything [52], which are trained on mixed dataset to enhance generality. Recent progress in generative methods, such as Marigold [28] and GeoWizard [15], demonstrates that the use of diffusion priors can facilitate zero-shot depth estimation with improved prediction details and smoothness. However, despite their impressive performance in complex scenarios, these single-image depth estimation methods often face challenges in maintaining consistency across multiple frames.
@@ 3.1.1 Diffusion Guided Sampling
$)\hat{\epsilon}_{\theta}(z_{t},t)=\epsilon_{\theta}(z_{t},t)+s(t)\cdot\nabla_{z_{t}}\mathcal{L}(c,f(\hat{z_{0}}$)) (2)
@@ 4 Experiments
Method Type Video ScanNet GMU KITCHEN Video KITTI Bonn Video ScanNet
@@
DepthAnythingV2 S 150 0.148 0.776 0.188 0.645 190 0.150 0.788 0.102 0.916 150 0.040
Marigold S 0.179 0.740 0.214 0.584 0.143 0.820 0.098 0.932 0.131
NVDS V 0.199 0.647 0.231 0.542 0.212 0.658 0.163 0.771 0.047
ChronoDepth V 0.172 0.749 0.196 0.650 0.178 0.733 0.092 0.932 0.028
DepthCrafter V 0.141 0.799 0.143 0.795 0.114 0.879 0.095 0.917 0.024
DepthSync(Ours) V 0.113 0.870 0.113 0.881 0.110 0.887 0.069 0.972 0.019
DepthAnythingV2 S 210 0.160 0.743 0.189 0.652 270 0.156 0.777 0.117 0.887 210 0.040
Marigold S 0.191 0.710 0.221 0.598 0.146 0.809 0.113 0.901 0.143
NVDS V 0.208 0.622 0.223 0.578 0.217 0.647 0.176 0.741 0.048
ChronoDepth V 0.182 0.726 0.201 0.659 0.185 0.707 0.112 0.898 0.028
DepthCrafter V 0.156 0.757 0.155 0.761 0.116 0.867 0.121 0.863 0.025
DepthSync(Ours) V 0.127 0.836 0.113 0.884 0.111 0.882 0.077 0.960 0.021
DepthAnythingV2 S 270 0.161 0.738 0.181 0.681 350 0.161 0.760 0.125 0.870 270 0.041
Marigold S 0.195 0.699 0.219 0.613 0.151 0.796 0.123 0.884 0.148
NVDS V 0.211 0.614 0.222 0.586 0.219 0.643 0.180 0.722 0.048
ChronoDepth V 0.184 0.716 0.201 0.659 0.185 0.705 0.124 0.875 0.029
DepthCrafter V 0.161 0.745 0.154 0.764 0.119 0.859 0.130 0.841 0.025
DepthSync(Ours) V 0.136 0.814 0.120 0.860 0.112 0.874 0.083 0.946 0.021
DepthAnythingV2 S 330 0.164 0.729 0.191 0.655 430 0.160 0.762 0.127 0.861 330 0.041
Marigold S 0.198 0.691 0.228 0.611 0.150 0.800 0.126 0.871 0.166
NVDS V 0.214 0.608 0.229 0.568 0.221 0.641 0.182 0.712 0.049
ChronoDepth V 0.187 0.709 0.217 0.621 0.186 0.706 0.130 0.855 0.029
DepthCrafter V 0.168 0.725 0.153 0.756 0.123 0.853 0.134 0.834 0.025
DepthSync(Ours) V 0.138 0.802 0.128 0.842 0.115 0.871 0.090 0.937 0.021
DepthAnythingV2 S 390 0.167 0.724 0.193 0.645 510 0.158 0.766 0.128 0.861 390 0.041
Marigold S 0.202 0.689 0.228 0.613 0.149 0.800 0.126 0.871 0.148
NVDS V 0.216 0.603 0.232 0.561 0.221 0.641 0.185 0.701 0.050
ChronoDepth V 0.190 0.704 0.222 0.619 0.186 0.705 0.132 0.851 0.030
DepthCrafter V 0.169 0.721 0.161 0.746 0.124 0.850 0.138 0.817 0.025
DepthSync(Ours) V 0.150 0.780 0.128 0.848 0.116 0.870 0.091 0.934 0.021
DepthAnythingV2 S 450 0.169 0.722 0.193 0.644 590 0.158 0.769 0.131 0.856 450 0.041
Marigold S 0.205 0.686 0.241 0.597 0.148 0.803 0.128 0.865 0.147
NVDS V 0.219 0.597 0.235 0.558 0.223 0.640 0.189 0.692 0.050
ChronoDepth V 0.193 0.699 0.225 0.612 0.193 0.691 0.132 0.853 0.030
DepthCrafter V 0.171 0.716 0.160 0.748 0.173 0.727 0.143 0.801 0.025
DepthSync(Ours) V 0.154 0.769 0.131 0.837 0.117 0.869 0.093 0.929 0.021
@@
[…]
@@ 4.2 Zero-Shot Depth Estimation Results
Figure 4: Qualitative Comparison for Long Video Depth Estimation. We compare predictions on 450-frame ScanNet videos and 590-frame Bonn videos. Slicing along the timeline at the green-line position (first column), we concatenate results to visualize temporal changes in color and depth. Red boxes highlight depth scale inconsistencies, while blue boxes mark depth inaccuracies in previous methods. Our method shows superior depth accuracy and scale consistency over time. Depth error map comparisons are provided in Appendix.
@@ 5 Conclusion
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 62495061, 62361146854), and Tsinghua-Tencent Joint Laboratory for Internet Innovation Technology.
@@ Appendix A Implementation Details
A.1 Hyperparameters
We employ the Euler scheduler [26] with 5 denoising steps, as per DepthCrafter [22], we apply both guidance terms sequentially in the last two steps: geometry guidance first to refine geometry within each window, followed by scale guidance to synchronize scales across windows.
Scale guidance uses forward guidance [1], back-propagating the MSE loss gradients (Eqn. (4) in the main paper) to the noise prediction. This process iterates 1000 times in the final denoising step. The base learning rate $\mathrm{lr}$ (Eqn. (2) in the main paper) is set to 6e4 for the second-to-last step and 1e6 for the final step, decaying as:
$s(t)=\mathrm{lr}\times 0.98^{\mathrm{iter}//100}$ (6)
Early termination occurs if the loss falls below 5e45e-4.
Geometry guidance uses backward guidance, optimizing the predicted depth decoded from the clean latent (Eqn. (1) in the main paper). We set the base learning rate to 0.01 for the second-to-last step and 0.02 for the final step, with 80 optimization iterations per step. We use the AdamW optimizer (betas = (0.9, 0.999), weight decay = 1e-4) and a cosine annealing scheduler (eta min = base lr * 0.01, T max = 80).
A.2 Loss Terms in Geometry Guidance
@@
A.2.1 Global Geometric Constraints constraints
@@
A.2.2 Local Geometric Constraints
@@ Appendix B Comparison with Reconstruction Methods
We compare our method with representative reconstruction methods on the first 10 ScanNet test scenes. COLMAP[41] and 2DGS [Huang2DGS2024] (uses COLMAP poses) fail on 3 scenes due to failed COLMAP pose estimation, while our method reconstructs all scenes successfully. For shared scenes, our approach achieves more accurate depth and poses (Table 4).
@@ Appendix C Inference Cost and Supplementary Evaluation
Dataset ScanNet GMU Kitchen KITTI Bonn
[…]
ChronoDepth [42] 0.172 0.749 0.196 0.650 0.178 0.733 0.092 0.932
DepthCrafter [22] 0.141 0.799 0.143 0.795 0.114 0.879 0.095 0.917
DUSt3R [47] 0.059 0.972 0.069 0.968 0.155 0.777 0.075 0.935
MonST3R [60] 0.067 0.959 0.087 0.920 0.192 0.692 0.046 0.975
Ours-S 0.128 0.834 0.126 0.844 0.112 0.882 0.070 0.968
Ours 0.113 0.870 0.113 0.881 0.110 0.887 0.069 0.972
ChronoDepth [42] 0.193 0.699 0.225 0.612 0.193 0.691 0.132 0.853
DepthCrafter [22] 0.171 0.716 0.160 0.748 0.173 0.727 0.143 0.801
DUSt3R [47] 0.090 0.934 0.127 0.828 0.174 0.731 0.115 0.889
MonST3R [60] 0.099 0.911 0.140 0.817 0.232 0.620 0.086 0.918
Ours-S 0.157 0.757 0.138 0.807 0.119 0.863 0.094 0.926
Ours 0.154 0.769 0.131 0.837 0.117 0.869 0.093 0.929
@@ Appendix E Supplementary Qualitative Examples
We provide additional qualitative examples to supplement the main paper. Figure 6, Figure 7, Figure 8 and Figure 9 demonstrate that our DepthSync achieves better scale synchronization between windows for long videos and more accurate geometric consistency within each window.
@@
Figure 7: Depth Error Map Comparison on Long Video Depth Consistency. Supplementary visualization of absolute error maps for qualitative analysis (see Figure 4 in the main paper and Figure 6). Blue indicates low error, while red corresponds to high error.